# AudioMoG: Guiding Audio Generation with Mixture-of-Guidance

### v1

@@ Abstract
Guidance methods have demonstrated significant improvements in cross-modal audio generation, including text-to-audio (T2A) and video-to-audio (V2A) generation. The popularly adopted method, classifier-free guidance (CFG), steers generation by emphasizing condition alignment, enhancing fidelity but often at the cost of diversity. Recently, autoguidance (AG) has been explored for audio generation, encouraging the sampling to faithfully reconstruct the target distribution and showing increased diversity. Despite these advances, they usually rely on a single guiding principle, e.g., condition alignment in CFG or score accuracy in AG, leaving the full potential of guidance for audio generation untapped. In this work, we explore enriching the composition of the guidance method and present a mixture-of-guidance framework, AudioMoG. Within the design space, AudioMoG can exploit the complementary advantages of distinctive guiding principles by fulfilling their cumulative benefits. With a reduced form, AudioMoG can consider parallel complements or recover a single guiding principle, without sacrificing generality. We experimentally show that, given the same inference speed, AudioMoG approach consistently outperforms single guidance in T2A generation across sampling steps, concurrently showing advantages in V2A, text-to-music, and image generation. These results highlight a “free lunch” in current cross-modal audio generation systems: higher quality can be achieved through mixed guiding principles at the sampling stage without sacrificing inference efficiency. Demo samples are available at: https://audio-mog.github.io.
1 Introduction
@@
Audio generation conditioned on text and video information, known as text-to-audio (T2A) and video-to-audio (V2A) generation, has witnessed significant advancements in recent studies. Typically, these systems generate an audio latent in a small space compressed from the audio waveform or the mel-spectrogram, indicated by the learned text embeddings (Kreuk et al., 2022; Liu et al., 2023; 2024a; Huang et al., 2023b; Evans et al., 2024; Jiang et al., 2025) or encoded video representations (Xu et al., 2024; Wang et al., 2024a; Du et al., 2023; Luo et al., 2023). Recent efforts have enhanced cross-modal audio generation quality through various perspectives, such as data augmentation (Huang et al., 2023b; a), condition information (Jeong et al., 2024; Wang et al., 2024b; Liu et al., 2023; Li et al., 2024a; Evans et al., 2024; Jiang et al., 2025), generative models (Kreuk et al., 2022; Liu et al., 2023; 2024a), network architecture (Huang et al., 2023a; Evans et al., 2024; Hung et al., 2024), and compression networks (Liu et al., 2023; Evans et al., 2024; 2025). However, most of these improvements require retraining the model from scratch or with significant overhead.
At the sampling stage, guidance methods have proven effective in enhancing the overall audio generation quality, where classifier-free guidance (CFG) (Ho & Salimans, 2022) is popularly adopted in modern cross-modal audio generation systems (Liu et al., 2023; 2024a; Cheng et al., 2025). By emphasizing the indication of condition signals, namely text or video, CFG can improve the audio generation results under an appropriate guidance scale, while it may therefore sacrifice generation diversity. Recently, autoguidance (AG) (Karras et al., 2024a) is proposed in class-conditioned image generation and has been extended to audio generation by ETTA (Lee et al., 2025). Different from CFG strengthening condition alignment, it guides the diffusion model with a weaker version, encouraging the generation process to faithfully reconstruct the target distribution.
@@
While these guidance methods are advantageous over unguided diffusion sampling, they usually strengthen the generation direction with a single guiding principle, e.g., condition alignment in CFG and score accuracy in AG. It remains underexplored if stronger generation results can be achieved by mixing distinctive guidance methods, while maintaining inference efficiency. In this work, we explore enriching the composition of the guidance method and present a mixture-of-guidance framework, named AudioMoG, to simultaneously consider distinctive guiding principles rather than depend solely on one of them. Firstly, we revisit the design of guidance strategies in audio generation, where we analyze the behaviors and limitations of the widely-used CFG and recent AG, respectively. We demonstrate that, CFG enhances synthesis quality through an entangled effect of score correction and condition alignment amplification, which complicates independent control over quality and diversity—particularly as improvements in the unconditional model diminish the correction signal. In comparison, AG employs a weaker conditional model to isolate the score correction effect, achieving more accurate score estimation to enable quality improvements, though its effectiveness can be sensitive to the choice of the weak model (Karras et al., 2024a; Lee et al., 2025).
Based on these insights, we demonstrate the mechanism of AudioMoG, an improved sampling framework that can fully exploit the complementary advantages of diverse guidance methods. Within the design space, AudioMoG can fulfill the cumulative advantage by progressively harnessing the strengths of diverse guiding principles, reaching the performance of further refining condition-aligned term empowered by CFG with AG, or strengthening both conditional and unconditional score estimation results with AG before CFG. As a degraded form, AudioMoG can fulfill parallel complements or ultimately recover the guidance method considering a single principle, such as CFG or AG, without sacrificing the generality as a mixture framework. Especially, in AudioMoG, we empirically observe that the bad version of the model can be trained using the same network architecture as the good version but with fewer iterations, or even taken directly from earlier checkpoints, avoiding the dedicated design proposed in AG (Karras et al., 2024a) or the sensitivity for audio generation mentioned in ETTA (Lee et al., 2025). Our contributions are summarized as follows.
  • We present a mixture-of-guidance framework for audio generation, achieving improved synthesis quality while maintaining the inference efficiency in comparison with the single guiding principle.
  • Within the design space, AudioMoG can fulfill the cumulative advantages of diverse guidance methods, progressively expressing their strengths, as well as allowing parallel complements or a single principle as a degraded form.
  • Experimental validation on diverse cross-modal audio generation and image generation tasks demonstrates that under the same inference speed, AudioMoG consistently outperforms both CFG and AG. Compared to CFG, we improve FAD from 1.76 to 1.38 in T2A (Evans et al., 2024), from 0.73 to 0.68 in V2A, and from 2.36 to 1.92 in text-to-music generation. Compared to AG, we improve FID from 1.60 to 1.47 in image generation (Karras et al., 2024b).
2 Preliminaries
In this section, we introduce the foundation of latent diffusion models and guided audio generation.
Diffusion-based audio generation system. In T2A and V2A generation systems, audio signals $\bm{x}$ are first compressed into a small latent space $\bm{z}$ with a compression network. Then, latent diffusion models are popularly adopted to learn the generation of audio latent from a simple prior distribution, e.g., the standard Gaussian noise distribution $\mathcal{N}(\bm{0},\bm{I})$, conditioned on the text prompt or video input. At the training stage, a forward process is introduced to transform the audio latent at $t=0$ into a noisy latent with
@@ 2 Preliminaries
where $\bm{c}$ is the text embedding or encoded video features to indicate audio generation. Given the well-trained noise predictor, a reverse process iteratively reconstructs the audio latent from the prior distribution with
$p_{\theta}(\bm{z}_{0:T-1}|\bm{z}_{T},\bm{c})=p(\bm{z}_{T})\prod_{t=1}^{T}p_{\theta}(\bm{z}_{t-1}|\bm{z}_{t},\bm{c}).$ (3)
In sampling, each reverse transition $p_{\theta}(\bm{z}_{s}|\bm{z}_{t},\bm{c})$ at the time steps $0\leq s<t\leq T$ follows a Gaussian distribution $\mathcal{N}(\bm{z}_{s},\mu_{s|t}(\bm{z}_{t},t,\bm{c}),\sigma_{s|t}^{2}\bm{I})$. The mean and variance are parameterized as
$\mu_{s|t}(\bm{z}_{t},t,\bm{c})=\sqrt{\bar{\alpha}_{s|t}}(\bm{z}_{t}-\frac{1-\bar{\alpha}_{t|s}}{\sqrt{1-\bar{\alpha}_{t}}}\bm{\epsilon_{\theta}}(\bm{z}_{t},t,\bm{c})),\quad\sigma^{2}_{s|t}=(1-\bar{\alpha}_{t|s})\frac{1-\bar{\alpha}_{s}}{1-\bar{\alpha}_{t}},$ (4)
@@
Classifier-Free Guidance. CFG (Ho & Salimans, 2022) is one of the most commonly used strategies in diffusion models for conditional generation. During training, the conditional signal $c$ is randomly replaced with the null condition $\mathcal{\varnothing}$ with a fixed probability $p_{\text{uncond}}$ (a.k.a., random label dropout), allowing the model to learn both conditional and unconditional noise predictors, $\bm{\epsilon_{\theta}}(\bm{z}_{t},t,c)$ and $\bm{\epsilon_{\theta}}(\bm{z}_{t},t)$. At inference time, CFG combines the two predictors as follows:
$\bm{\epsilon}_{\text{CFG}}(\bm{z}_{t},t,\bm{c})=\bm{\epsilon}_{\theta}(\bm{z}_{t},t)+w\left(\bm{\epsilon}_{\theta}(\bm{z}_{t},t,\bm{c})-\bm{\epsilon}_{\theta}(\bm{z}_{t},t)\right),$ (5)
where $w\geq 1$ denotes the guidance scale that adjusts the strength of the conditional signal. When $w=1$, CFG recovers the conditional diffusion model $\bm{\epsilon}_{\theta}(\bm{z}_{t},t,\bm{c})$. Larger values of $w$ encourage samples to align more closely with the conditioning signal, potentially enhancing generation quality.
@@
Figure 2: Illustration of guidance methods on the fractal-like 2D distribution from Karras et al. (2024a). (a) Ground truth distribution (orange class). (b) Unguided conditional sampling generates outliers. (c) Classifier-Free Guidance ($w=3$) with a well-trained unconditional model struggles to remove outliers. (d) Autoguidance ($w=3$) improves score estimation and removes outliers without reducing diversity. (e) Parallel Guidance exhibits mode dropping similar to CFG. (f) Hierarchical Guidance eliminates outliers and provides more controllable condition alignment.
@@
Autoguidance. AG (Karras et al., 2024a) proposes guiding a diffusion model using a weaker version of itself:
$\bm{\epsilon}_{\text{AG}}(\bm{z}_{t},t,\bm{c})=\bm{\epsilon}_{\theta_{\text{bad}}}(\bm{z}_{t},t,\bm{c})+w\left(\bm{\epsilon}_{\theta}(\bm{z}_{t},t,\bm{c})-\bm{\epsilon}_{\theta_{\text{bad}}}(\bm{z}_{t},t,\bm{c})\right),$ (6)
@@
where $\theta_{\text{bad}}$ refers to the weak model with smaller size or less training, and $\bm{c}$ can be replaced with $\varnothing$. The underlying motivation stems from the observation that the score-matching objective in diffusion models promotes mode coverage, often leading to noisy or inaccurate estimates. By contrastive amplification of the difference between a strong and weak model, AG seeks to improve the score estimation quality. Following a similar rationale, recent works (Kasymov et al., 2024; Phunyaphibarn et al., 2025; Zhong et al., 2025) guide the fine-tuned model using the pre-fine-tuned one. While the specific formulations differ, the core principle remains consistent: leveraging the weak-strong discrepancy to guide improvement.
3 AudioMoG
3.1 Analysis
@@
2D toy example. We first provide an analysis of the CFG and AG methods to illustrate their respective guiding principle. We adopt a 2D toy example introduced in Karras et al. (2024a), where a small denoiser is trained on synthetic data to learn conditional diffusion. The 2D dataset is designed to exhibit low local dimensionality, characterized by highly anisotropic and narrow support, as well as a hierarchical emergence of local detail as shown in Figure 2(a), mimicking real-world data manifolds (Karras et al., 2024a). As shown in Figure 2(b), the denoiser network learns a suboptimal score function, leading to scattered and unlikely outliers under unguided generation, i.e., conditional diffusion sampling. Additional details on the experimental setup are provided in Appendix E.
@@
CFG effects. For guided generation results, CFG improves sample quality by contrasting the conditional and unconditional models. The unconditional model is arguably relatively under-trained due to the inherent difficulty of the unconditional task and the low label dropout rate (e.g., 10% for audio generation in Liu et al. (2023); Deepanway et al. (2023); Evans et al. (2024)). Consequently, the CFG effect in Equation 5 is mixed, as formulated in the following guidance decomposition:
$\nabla_{\bm{x}}\log p(\bm{x}|\bm{c})-\nabla_{\bm{x}}\log p(\bm{x}|{\varnothing})=[\nabla_{\bm{x}}\log\mathbb{E}_{c}p(\bm{x}|\bm{c})-\nabla_{\bm{x}}\log p(\bm{x}|{\varnothing})]+\nabla_{\bm{x}}\log p(\bm{c}|\bm{x}).$ (7)
@@
Namely, with an under-trained unconditional model, the CFG direction entangles two components: (1) score correction from weak-strong contrast that eliminates dispersed outliers, and (2) condition alignment amplification that may skew the distribution and reduce diversity. The quality improvement observed with CFG can be ideally attributed to the first factor. However, the entanglement makes it difficult for CFG to independently control diversity and quality, as a better unconditional model yields weaker signals for score correction. This effect is visualized in Figure 2(c), where CFG with a sufficiently trained unconditional model fails to eliminate dispersion.
@@
AG effects. On the other hand, AG avoids this entanglement by employing a weaker but still conditional version of the model, thus isolating the improvement direction without losing diversity as shown in Figure 2(d). However, its effectiveness hinges on the availability of a suitably degraded weak model. Constructing such a model in practice can be nontrivial (Karras et al., 2024a; Lee et al., 2025; Jeon, 2025; Hyung et al., 2025), especially when model degradation does not align well with real score estimation errors and the weak-strong contrast cannot provide meaningful directions. In such cases, incorporating additional sources of guidance may be necessary to achieve more robust quality gains. For example, when training data quality is suboptimal, CFG often yields sharper and more prompt-consistent generations due to its “lower-temperature” behavior (Bradley & Nakkiran, 2024), creating a more favorable distribution that can complement AG.

### v2

@@ Abstract
The design of diffusion-based audio generation systems has been investigated from diverse perspectives, such as data space, network architecture, and conditioning techniques, while most of these innovations require model re-training. In sampling, classifier-free guidance (CFG) has been uniformly adopted to enhance generation quality by strengthening condition alignment. However, CFG often compromises diversity, resulting in suboptimal performance. Although the recent autoguidance (AG) method proposes another direction of guidance that maintains diversity, its direct application in audio generation has so far underperformed CFG. In this work, we introduce AudioMoG, an improved sampling method that enhances text-to-audio (T2A) and video-to-audio (V2A) generation quality without requiring extensive training resources. We start with an analysis of both CFG and AG, examining their respective advantages and limitations for guiding diffusion models. Building upon our insights, we introduce a mixture-of-guidance framework that integrates diverse guidance signals with their interaction terms (e.g., the unconditional bad version of the model) to maximize cumulative advantages. Experiments show that, given the same inference speed, our approach consistently outperforms single guidance in T2A generation across sampling steps, concurrently showing advantages in V2A, text-to-music, and image generation. Demo samples are available at: audiomog.github.io.
I Introduction
@@
Audio generation conditioned on text and video information, known as text-to-audio (T2A) and video-to-audio (V2A) generation, have witnessed significant advancements in recent studies. Typically, these systems generate an audio latent in a small space compressed from the audio waveform or the mel-spectrogram, indicated by text embeddings [35, 42, 43, 23, 13] or encoded video representations [59, 55, 12, 46]. Recent efforts have enhanced cross-modal audio generation quality through various perspectives, such as data augmentation [23, 22], condition information [27, 56, 42, 39, 13], generative models [35, 42, 43], network architecture [22, 13, 24], and compression networks [42, 13, 14]. However, most of these improvements require retraining the model from scratch or with significant overhead.
At the sampling stage, classifier-free guidance (CFG) is the dominant approach for improving generation quality by amplifying the effect of text conditions, while its performance remains unsatisfactory due to reduced diversity. Recently, alternative guidance mechanisms such as autoguidance (AG) [31], which guides a diffusion model with a weak version of itself, have shown promise to outperform CFG in image [31, 62], motion [26], and video [25] generation. However, AG has not proven effective in audio generation settings. For instance, ETTA [38] investigates AG for T2A generation and finds that although it promotes output diversity, it is sensitive to the choice of weak model and fails to match CFG in terms of objective metrics. The potential and proper usage of guidance in cross-modal audio generation remains underexplored for sampling optimization.
In this work, we first revisit the design of guidance strategies in audio generation, where we analyze the behaviors and limitations of the widely-used CFG and recent AG, respectively. We demonstrate that, CFG enhances synthesis quality through an entangled effect of score correction and condition alignment amplification, which complicates independent control over quality and diversity—particularly as improvements in the unconditional model diminish the correction signal. In contrast, AG employs a weaker conditional model to isolate the score correction effect, enabling quality improvements without compromising diversity, though its effectiveness can be sensitive to the choice of the weak model.
Based on these insights, we present AudioMoG, a mixture-of-guidance sampling framework towards enhancing the generation quality without requiring heavy retraining resources, which fully exploits the complementary advantages of diverse guidance methods. Specifically, AudioMoG presents two mixture strategies: Parallel Guidance and Hierarchical Guidance, where diverse guidance methods are integrated in parallel or hierarchically. The first strategy concurrently leverages the benefits of CFG and AG, while it can suffer from suboptimal results due to the potential conflicts behind the behaviors of different guidance methods. With the second strategy, we enable a more structured integration, progressively harnessing their strengths to yield cumulative advantages. For example, CFG can first emphasize condition alignment for both the strong and weak models, and then AG further strengthens the generation quality by reducing the score estimation errors between CFG-aligned terms. In AudioMoG, we observe that the bad version of the model can be trained using the same network architecture as the good one but with fewer iterations, or even taken directly from earlier checkpoints, avoiding the dedicated design [31] or the sensitivity [38] mentioned in previous works.
Our contributions are summarized as follows.
  • We make the first attempt to enhance the guidance method of audio generation and improve the generation quality in a computationally lightweight or even training-free manner.
  • We propose a novel guidance method, AudioMoG, where we present a mixture-of-guidance framework to maximize the cumulative advantages of different guidance methods.
  • Experimental validation on diverse cross-modal audio generation and image generation tasks demonstrates that under the same inference speed, AudioMoG consistently outperforms both CFG and AG. Compared to CFG, we improve FAD from 1.76 to 1.38 in T2A [13], from 0.73 to 0.68 in V2A, and from 2.36 to 1.92 in text-to-music generation. Compared to AG, we improve FID from 1.60 to 1.47 in image generation [30].
@@
II Diffusion-based audio generation system
In T2A and V2A generation systems, audio signals $\bm{x}$ are first compressed into a small latent space $\bm{z}$ with a compression network. Then, latent diffusion models are popularly adopted to learn the generation of audio latent from a simple prior distribution, e.g., the standard Gaussian noise distribution $\mathcal{N}(\bm{0},\bm{I})$, conditioned on the text prompt or video input. At the training stage, a forward process is introduced to transform the audio latent at $t=0$ into a noisy latent with
@@ II Diffusion-based audio generation system
where $\bm{c}$ is the text embedding or encoded video features to indicate audio generation. In sampling, each reverse transition $p_{\theta}(\bm{z}_{s}|\bm{z}_{t},\bm{c})$ at the time steps $0\leq s<t\leq T$ follows a Gaussian distribution $\mathcal{N}(\bm{z}_{s};\mu_{s|t}(\bm{z}_{t},t,\bm{c}),\sigma_{s|t}^{2}\bm{I})$. The mean and variance are parameterized as
$\mu_{s|t}(\bm{z}_{t},t,\bm{c})=\sqrt{\bar{\alpha}_{s|t}}(\bm{z}_{t}-\frac{1-\bar{\alpha}_{t|s}}{\sqrt{1-\bar{\alpha}_{t}}}\bm{\epsilon_{\theta}}(\bm{z}_{t},t,\bm{c})),$ (3)
$\sigma^{2}_{s|t}=(1-\bar{\alpha}_{t|s})\frac{1-\bar{\alpha}_{s}}{1-\bar{\alpha}_{t}},$ (4)
@@
III Guidance methods
We begin by introducing two representative guidance methods for diffusion models: the widely adopted Classifier-Free Guidance (CFG) and the recently proposed Autoguidance (AG). We then analyze their underlying mechanisms and compare their respective advantages and limitations.
Classifier-Free Guidance
CFG [20] is one of the most commonly used strategies in diffusion models for conditional generation. During training, the conditional signal $c$ is randomly replaced with the null condition $\mathcal{\varnothing}$ with a fixed probability $p_{\text{uncond}}$ (aka., random label dropout), allowing the model to learn both conditional and unconditional noise predictors, $\bm{\epsilon_{\theta}}(\bm{z}_{t};t,c)$ and $\bm{\epsilon_{\theta}}(\bm{z}_{t};t)$. At inference time, CFG combines the two predictors as follows:
$\bm{\epsilon}_{\text{CFG}}(\bm{z}_{t};t,\bm{c})=\bm{\epsilon}_{\theta}(\bm{z}_{t};t)+w\left(\bm{\epsilon}_{\theta}(\bm{z}_{t};t,\bm{c})-\bm{\epsilon}_{\theta}(\bm{z}_{t};t)\right),$ (5)
where $w\geq 1$ denotes the guidance scale that adjusts the strength of the conditional signal. Larger values of $w$ encourage samples to align more closely with the conditioning signal, potentially enhancing generation quality.
@@ III Guidance methods
Figure 2: Illustration of guidance methods on the fractal-like 2D distribution from [31]. (a) Ground truth distribution (orange class). (b) Unguided conditional sampling generates outliers. (c) CFG ($w=3$) with a well-trained unconditional model struggles to remove outliers. (d) AG ($w=3$) improves score estimation and removes outliers without reducing diversity. (e) Parallel Guidance exhibits mode dropping similar to CFG. (f) Hierarchical Guidance eliminates outliers and provides more controllable condition alignment.
Autoguidance
@@
AG [31] proposes guiding a diffusion model using a weaker version of itself:
$\bm{\epsilon}_{\text{AG}}(\bm{z}_{t};t,\bm{c})=\bm{\epsilon}_{\theta_{\text{bad}}}(\bm{z}_{t};t,\bm{c})+w\left(\bm{\epsilon}_{\theta}(\bm{z}_{t};t,\bm{c})-\bm{\epsilon}_{\theta_{\text{bad}}}(\bm{z}_{t};t,\bm{c})\right),$ (6)
@@
where $\theta_{\text{bad}}$ refers to the weak model with smaller size or less training, and $\bm{c}$ can be replaced with $\varnothing$. The underlying motivation stems from the observation that the score-matching objective in diffusion models promotes mode coverage, often leading to noisy or inaccurate estimates. By contrastive amplification of the difference between a strong and weak model, AG seeks to improve the score estimation quality. Following a similar rationale, recent works [32, 50, 63] leverage the pre-fine-tuned model to guide the model after fine-tuning. While the specific formulations differ, the core principle remains consistent: leveraging the weak-strong discrepancy to guide improvement.
Analysis
To demonstrate the mechanisms of CFG and AG, we adopt a 2D toy example introduced in [31], where a small denoiser is trained on synthetic data to learn conditional diffusion. Additional details on the experimental setup are provided in Appendix E.
@@
The 2D dataset is designed to exhibit low local dimensionality, characterized by highly anisotropic and narrow support, as well as a hierarchical emergence of local detail (Fig. 2(a)), mimicking real-world data manifolds [31]. Due to limited capacity, the small denoiser network learns a suboptimal score function, leading to scattered and unlikely outliers under unguided sampling (Fig. 2(b)).
@@
CFG improves sample quality by contrasting the conditional and unconditional models. The unconditional model is arguably relatively under-trained due to the inherent difficulty of the unconditional task and the low label dropout rate (usually 10%). Consequently, the CFG effect in (5) is mixed, as formulated in the following score decomposition:
$\nabla_{\bm{x}}\log p(\bm{x}|\bm{c})-\nabla_{\bm{x}}\log p(\bm{x}|\varnothing)=\bigl[\nabla_{\bm{x}}\log\mathbb{E}_{c}\,p(\bm{x}|\bm{c})$
$-\nabla_{\bm{x}}\log p(\bm{x}|\varnothing)\bigr]+\nabla_{\bm{x}}\log p(\bm{c}|\bm{x}).$ (7)
@@
With an under-trained unconditional model, the CFG direction entangles two components: (1) score correction from weak-strong contrast that eliminates dispersed outliers, and (2) condition alignment amplification that may skew the distribution and reduce diversity. The quality improvement observed with CFG can be ideally attributed to the first factor. However, the entanglement makes CFG difficult to independently control diversity and quality, as a better unconditional model yields weaker signals for score correction. This effect is visualized in Fig. 2(c), where CFG with a sufficiently trained unconditional model fails to eliminate dispersion.
@@
AG, on the other hand, avoids this entanglement by employing a weaker but still conditional version of the model, thus isolating the improvement direction without losing diversity (Fig. 2(d)). However, its effectiveness hinges on the availability of a suitably degraded weak model. Constructing such a model in practice can be nontrivial, especially when model degradation does not align well with real score estimation errors and the weak-strong contrast cannot provide meaningful directions. In such cases, incorporating additional sources of guidance may be necessary to achieve more robust quality gains. For example, when training data quality is suboptimal, CFG often yields sharper and more prompt-consistent generations due to its “lower-temperature” behavior [3], creating a more favorable distribution that can complement AG.
IV AudioMoG